# SLOT: Sample-specific Language Model Optimization at Test-time

### v1

@@ Abstract
We propose SLOT (Sample-specific Language Model Optimization at Test-time), a novel and parameter-efficient test-time inference approach that enhances a language model’s ability to more accurately respond to individual prompts. Existing Large Language Models (LLMs) often struggle with complex instructions, leading to poor performances on those not well represented among general samples. To address this, SLOT conducts few optimization steps at test-time to update a light-weight sample-specific parameter vector. It is added to the final hidden layer before the output head, and enables efficient adaptation by caching the last layer features during per-sample optimization. By minimizing the cross-entropy loss on the input prompt only, SLOT helps the model better aligned with and follow each given instruction. In experiments, we demonstrate that our method outperforms the compared models across multiple benchmarks and LLMs. For example, Qwen2.5-7B with SLOT achieves an accuracy gain of 8.6% on GSM8K from 57.54% to 66.19%, while DeepSeek-R1-Distill-Llama-70B with SLOT achieves a SOTA accuracy of 68.69% on GPQA among 70B-level models. Our code is available at https://github.com/maple-research-lab/SLOT.
@@ 1 Introduction
Large Language Models (LLMs) brown2020language have shown strong general capabilities in text generation, comprehension, and interaction. To further boost their performances, test-time scaling has been proposed as a strategy that allocates additional computation during inference to generate more accurate responses for each individual question. In this context, a distinct paradigm known as Test-Time Adaptation (TTA) has emerged, which focuses on optimizing model parameters for individual prompts. Early approaches sun2020test implemented instance-specific online adaptation via self-supervision, leveraging loss functions such as entropy loss or reconstruction loss to guide weight updates. More recently, Test-Time Reinforcement Learning (TTRL) li2025ttrl introduced the use of majority voting results as a reward signal during inference, enabling the model to adapt itself iteratively. However, TTA methods often suffer from high computational overhead due to the need for per-instance updates on large-scale models howard2024surprisingeffectiveness , and designing effective supervision signals for complex LLM tasks remains a significant challenge.
@@
We conduct experiments on a wide range of LLMs and benchmarks to validate the effectiveness of the SLOT. Particularly, SLOT boosts Qwen2.5-7B’s performance on GSM8K by an 8.6% gain (from 57.54% to 66.19%) and enables DeepSeek-R1-Distill-Llama-70B to achieve an accuracy of 68.69% on GPQA, a record 3.03% improvement for 70B opensource models. As shown in Figure 1, SLOT can successfully improve in both answer accuracy and format accuracy as more optimization steps are adopted.
@@ 2 Related work
In-Context Learning (ICL) and Chain of Thoughts (CoT) Large language models exhibit ICL capabilities brown2020language , where they adapt their behavior based on examples provided within the input prompt, without any gradient updates. The mechanisms underlying ICL are still under investigation min2022rethinking . ICL relies solely on the model’s forward pass and attention mechanisms operating over the context, while SLOT employs explicit, gradient-based optimization to modify the model’s internal state representation for the current input.
@@
Parameter-Efficient Fine-Tuning (PEFT) PEFT adapts pretrained models for downstream tasks during a training or fine-tuning phase by updating only a small number of parameters, keeping the bulk of the LLM frozen. Examples include Adapter Tuning houlsby2019parameter , LoRA hu2022lora , Prompt Tuning lester2021power , Prefix Tuning li2021prefix , and P-Tuning v2 liu2022ptuningv2 . Crucially, all these methods learn a single set of efficient parameters for an entire task or dataset during a dedicated fine-tuning phase. In contrast, SLOT operates exclusively during the inference phase and optimizes a temporary, sample-specific parameter $\delta$ that is discarded after processing the current input. SLOT does not aim for task-level adaptation but for enhancing the model’s processing of the specific instance being evaluated.
@@
Test-Time Adaptation (TTA) TTA aims to address distribution shifts encountered during deployment by dynamically adjusting pre-trained models using only unlabeled test instances, typically in a source-free manner wang2023comprehensivesurveytesttime ; gandelsman2024beyondmodeladaptation . Foundational TTA approaches, often originating in computer vision, include minimizing prediction uncertainty (exemplified by TENT wang2021tent which adapts Batch Normalization layers via entropy minimization) and Test-Time Training (TTT) using self-supervision (pioneered by Sun et al. sun2020testtimetraining with tasks like rotation prediction, and later extended to methods like masked autoencoding you2022testtimetrainingmasked and normalizing flows yu2022tttflow ). While these techniques and associated strategies like combined self-training zhang2023towardsrealworld or meta-learning for adaptation wang2023learningtolearn have seen broad application zhang2024stta ; wang2023augmentedtesttime ; darestani2022testtimetrainingcompressed , their principles are increasingly critical and actively explored for Large Language Models (LLMs). TTA often involves instance-specific online adaptation such as leveraging in-context learning for few-shot reasoning howard2024surprisingeffectiveness , adapting via retrieved data or input rewriting chen2024improvingblackbox , or employing Test-Time Reinforcement Learning li2025ttrl . However, applying TTA to massive LLMs exacerbates general challenges like error accumulation and catastrophic forgetting liu2021tttplusplus ; goyal2022onpitfalls , and introduces acute issues such as prohibitive computational overhead for per-instance updates howard2024surprisingeffectiveness , the difficulty of designing effective self-supervision for complex linguistic tasks, critical reliance on adaptation data quality howard2024surprisingeffectiveness ; chen2024improvingblackbox , and risks to model alignment and fairness. Consequently, developing robust, efficient, lightweight, and safe TTA methods, particularly tailored for LLMs, remains a key ongoing research direction.
@@
Test-Time Scaling (TTS) TTS aims to enhance LLMs’ performance, especially on complex reasoning tasks, by allocating additional computational resources during inference li2025ttrl , inspired by the benefits of increased ‘thinking time’ processing and often surveyed in works like li2025stopoverthinking ; liu2025surveyonllmtesttime ; zhang2025testtimecompute . This typically involves external TTS with fixed pre-trained models zhao2025can1bllm , leveraging inference-time techniques such as sampling-based methods like Self-Consistency (which aggregates multiple Chain-of-Thought paths wang2023selfconsistency ) and Best-of-N sampling (often guided by Reward Models zhao2025can1bllm ; albalak2025benchmarkingreasoning ), or more structured search-based approaches like Tree of Thoughts (ToT) that enable deliberate exploration and backtracking zhang2025testtimecompute ; yao2023treeofthoughts . More advanced strategies include adaptive Test-Time Compute, where models dynamically allocate resources based on perceived task difficulty or confidence madaan2024adaptiveinferencetime ; zhang2024learninghowhard . A critical insight from TTS research is that strategic inference-time computation can enable smaller LLMs to surpass larger counterparts zhao2025can1bllm , challenging traditional scaling paradigms focused solely on model size. However, TTS methods generally incur significant latency and computational costs, and their efficacy depends heavily on well-calibrated guidance mechanisms, making the balance between performance gains and practical constraints a key area of ongoing investigation.
@@ 3.2 Method
where $p(x_{i+1}|x_{1:i},\delta)$ is the output probability of the next token $x_{i+1}$ given the context $x_{1:i}$ and the current $\delta$, using the adapted hidden features $H^{\prime}=H+\delta$. The optimization can be performed with a standard gradient descent optimizer like AdamW loshchilov2019decoupledweightdecayregularization :
@@ 4.2 Models and benchmarks
  • GPQA: A graduate-level Google-Proof Q&A benchmark (rein2024gpqa, ) requiring deep domain knowledge and complex reasoning, primarily in STEM fields.
@@
  • C-Eval: A comprehensive Chinese language evaluation suite covering multiple subjects (huang2023cevalmultilevelmultidisciplinechinese, ). We report performance across its main categories (STEM, Social Science, Humanities, Other), its ’Hard’ subset, and the overall average.
@@ 4.4 SLOT for Reasoning
As reported in Table 2, SLOT increases the DeepSeek-R1-Distill-Llama-70B baseline on AIME24 from 63.33% to 73.33%, and GPQA from 65.66% to 68.69%, achieving the SOTA performance for this task with the open-source 70B baseline model.
@@
Model AIME24 Math500 GPQA
@@ 5 Conclusion
We introduced SLOT, a novel test-time adaptation approach that optimizes a lightweight, sample-specific parameter vector added to the final hidden features of large language model. By minimizing the cross-entropy loss on the input prompt itself over a few optimization steps, SLOT efficiently aligns the model more closely with the given instruction, leveraging cached last-layer features to ensure minimal computational overhead. Our experiments demonstrate significant performance improvements on challenging reasoning benchmarks such as GSM8K, AIME24, and GPQA across various LLMs; for instance, Qwen2.5-7B gained 8.65% on GSM8K, and DeepSeek-R1-Distill-Llama-70B achieved state-of-the-art GPQA results for its class. Analysis of the induced logit modulation suggests SLOT encourages deeper reasoning by enhancing the probability of relevant tokens. While SLOT offers a practical method for on-the-fly adaptation, future work might explore adaptive optimization schemes or its efficacy across an even broader range of tasks and modalities.

### v2

@@ Abstract
We propose SLOT (Sample-specific Language Model Optimization at Test-time), a novel and parameter-efficient test-time inference approach that enhances a language model’s ability to more accurately respond to individual prompts. Existing Large Language Models (LLMs) often struggle with complex instructions, leading to poor performances on those not well represented among general samples. To address this, SLOT conducts few optimization steps at test-time to update a light-weight sample-specific parameter vector. It is added to the final hidden layer before the output head, and enables efficient adaptation by caching the last layer features during per-sample optimization. By minimizing the cross-entropy loss on the input prompt only, SLOT helps the model better aligned with and follow each given instruction. In experiments, we demonstrate that our method outperforms the compared models across multiple benchmarks and LLMs. For example, Qwen2.5-7B with SLOT achieves an accuracy gain of 8.6% on GSM8K from 57.54% to 66.19%, while DeepSeek-R1-Distill-Llama-70B with SLOT achieves a SOTA accuracy of 68.69% on GPQA Diamond among 70B-level models. Our code is available at https://github.com/maple-research-lab/SLOT.
@@ 1 Introduction
Large Language Models (LLMs) mann2020language have shown strong general capabilities in text generation, comprehension, and interaction. To further boost their performances, test-time scaling has been proposed as a strategy that allocates additional computation during inference to generate more accurate responses for each individual question. In this context, a distinct paradigm known as Test-Time Adaptation (TTA) has emerged, which focuses on optimizing model parameters for individual prompts. Early approaches sun2020test implemented instance-specific online adaptation via self-supervision, leveraging loss functions such as entropy loss or reconstruction loss to guide weight updates. More recently, Test-Time Reinforcement Learning (TTRL) zuo2025ttrl introduced the use of majority voting results as a reward signal during inference, enabling the model to adapt itself iteratively. However, TTA methods often suffer from high computational overhead due to the need for per-instance updates on large-scale models akyurek2024surprising , and designing effective supervision signals for complex LLM tasks remains a significant challenge.
@@
We conduct experiments on a wide range of LLMs and benchmarks to validate the effectiveness of the SLOT. Particularly, SLOT boosts Qwen2.5-7B’s performance on GSM8K by an 8.6% gain (from 57.54% to 66.19%) and enables DeepSeek-R1-Distill-Llama-70B to achieve an accuracy of 68.69% on GPQA Diamond, a record 3.03% improvement for 70B opensource models. As shown in Figure 1, SLOT can successfully improve in both answer accuracy and format accuracy as more optimization steps are adopted.
@@ 2 Related work
In-Context Learning (ICL) and Chain of Thoughts (CoT) Large language models exhibit ICL capabilities mann2020language , where they adapt their behavior based on examples provided within the input prompt, without any gradient updates. The mechanisms underlying ICL are still under investigation min2022rethinkingroledemonstrationsmakes . ICL relies solely on the model’s forward pass and attention mechanisms operating over the context, while SLOT employs explicit, gradient-based optimization to modify the model’s internal state representation for the current input.
@@
Parameter-Efficient Fine-Tuning (PEFT) PEFT adapts pretrained models for downstream tasks during a training or fine-tuning phase by updating only a small number of parameters, keeping the bulk of the LLM frozen. Examples include Adapter Tuning houlsby2019parameter , LoRA hu2022lora , Prompt Tuning lester2021power , Prefix Tuning li2021prefix , and P-Tuning v2 liu2021p . Crucially, all these methods learn a single set of efficient parameters for an entire task or dataset during a dedicated fine-tuning phase. In contrast, SLOT operates exclusively during the inference phase and optimizes a temporary, sample-specific parameter $\delta$ that is discarded after processing the current input. SLOT does not aim for task-level adaptation but for enhancing the model’s processing of the specific instance being evaluated.
@@
Test-Time Adaptation (TTA) TTA aims to address distribution shifts encountered during deployment by dynamically adjusting pre-trained models using only unlabeled test instances, typically in a source-free manner xiao2024beyond . Foundational TTA approaches, often originating in computer vision, include minimizing prediction uncertainty (exemplified by TENT wang2020tent which adapts Batch Normalization layers via entropy minimization) and Test-Time Training (TTT) using self-supervision sun2020test with tasks like rotation prediction, and later extended to methods like masked autoencoding gandelsman2022test and normalizing flows osowiechi2023tttflow ). While these techniques and associated strategies like combined self-training su2024towards or meta-learning for adaptation sun2023learning have seen broad application xiong2023stta ; darestani2022test , their principles are increasingly critical and actively explored for Large Language Models (LLMs). TTA often involves instance-specific online adaptation such as leveraging in-context learning for few-shot reasoning akyurek2024surprising , adapting via retrieved data or input rewriting o2024improving , or employing Test-Time Reinforcement Learning zuo2025ttrl . However, applying TTA to massive LLMs exacerbates general challenges like error accumulation and catastrophic forgetting liu2021ttt++ ; zhao2023pitfalls , and introduces acute issues such as prohibitive computational overhead for per-instance updates akyurek2024surprising , the difficulty of designing effective self-supervision for complex linguistic tasks, critical reliance on adaptation data quality and risks to model alignment and fairness. Consequently, developing robust, efficient, lightweight, and safe TTA methods, particularly tailored for LLMs, remains a key ongoing research direction.
@@
Test-Time Scaling (TTS) TTS aims to enhance LLMs’ performance, especially on complex reasoning tasks, by allocating additional computational resources during inference, inspired by the benefits of increased ‘thinking time’ processing and often surveyed in works like sui2025stop ; li2025survey ; ji2025test . This typically involves external TTS with fixed pre-trained models liu2025can , leveraging inference-time techniques such as sampling-based methods like Self-Consistency (which aggregates multiple Chain-of-Thought paths wang2022self ) and Best-of-N sampling (often guided by Reward Models yu2025benchmarking ), or more structured search-based approaches like Tree of Thoughts (ToT) that enable deliberate exploration and backtracking yao2023tree . More advanced strategies include adaptive Test-Time Compute, where models dynamically allocate resources based on perceived task difficulty or confidence manvi2024adaptive ; damani2024learning . A critical insight from TTS research is that strategic inference-time computation can enable smaller LLMs to surpass larger counterparts liu2025can , challenging traditional scaling paradigms focused solely on model size. However, TTS methods generally incur significant latency and computational costs, and their efficacy depends heavily on well-calibrated guidance mechanisms, making the balance between performance gains and practical constraints a key area of ongoing investigation.
@@ 3.2 Method
where $p(x_{i+1}|x_{1:i},\delta)$ is the output probability of the next token $x_{i+1}$ given the context $x_{1:i}$ and the current $\delta$, using the adapted hidden features $H^{\prime}=H+\delta$. The optimization can be performed with a standard gradient descent optimizer like AdamW loshchilov2017decoupled :
@@ 4.2 Models and benchmarks
  • GPQA Diamond: A graduate-level Google-Proof Q&A benchmark (rein2024gpqa, ) requiring deep domain knowledge and complex reasoning, primarily in STEM fields.
@@
  • C-Eval: A comprehensive Chinese language evaluation suite covering multiple subjects (huang2023c, ). We report performance across its main categories (STEM, Social Science, Humanities, Other), its ’Hard’ subset, and the overall average.
@@ 4.4 SLOT for Reasoning
As reported in Table 2, SLOT increases the DeepSeek-R1-Distill-Llama-70B baseline on AIME24 from 63.33% to 73.33%, and GPQA Diamond from 65.66% to 68.69%, achieving the SOTA performance for this task with the open-source 70B baseline model.
@@
Model AIME24 Math500 GPQA Diamond
@@ 5 Conclusion
We introduced SLOT, a novel test-time adaptation approach that optimizes a lightweight, sample-specific parameter vector added to the final hidden features of large language model. By minimizing the cross-entropy loss on the input prompt itself over a few optimization steps, SLOT efficiently aligns the model more closely with the given instruction, leveraging cached last-layer features to ensure minimal computational overhead. Our experiments demonstrate significant performance improvements on challenging reasoning benchmarks such as GSM8K, AIME24, and GPQA Diamond across various LLMs; for instance, Qwen2.5-7B gained 8.65% on GSM8K, and DeepSeek-R1-Distill-Llama-70B achieved state-of-the-art GPQA Diamond results for its class. Analysis of the induced logit modulation suggests SLOT encourages deeper reasoning by enhancing the probability of relevant tokens. While SLOT offers a practical method for on-the-fly adaptation, future work might explore adaptive optimization schemes or its efficacy across an even broader range of tasks and modalities.